Evaluate the following function or inverse; then write the ordered pair f(x)=3x+1 and f^-1=x-1/3
f(0)

Answer by stanbon(75887)

Evaluate the following function or inverse; then write the ordered pair
f(x)=3x+1 and f^-1 = (x-1)/3

f(0) = 3*0+1 = 1
===================
Cheers,
Stan H.